Suppose you purchase asset A intending to keep it for 10 years (receiving 10 total dividend payments), at which time you will sell it for the same price you purchased it for. What is the present value of the returns from that purchase?

Suppose you receive asset B for free and plan to keep it for 10 years and then sell it, at which point its value will have risen to $1,000.00. What is the present value of the return from that asset? 


1
Expert's answer
2021-11-14T17:34:16-0500

Solution:

Assume the purchase price for asset A is $10,000.

Annual dividend payments = $2000

You will sell the asset at the same price you purchased at the end of year 5 = 2,000 "\times" 5 = $10,000

 

Present value of the returns:

Dividend returns for 10 yrs = 2,000 "\times" 10 = $20,000

Interest rate = 5"\%"

PV="\frac{FV}{(1 + i)^{n} }=\frac{20,000}{(1 + 0.05)^{10} } = \frac{20,000}{1.62889} = \$12,278.30"


Assuming annual returns = $200

Its value will have risen to $1000 after 5 yrs

Total returns = 200 "\times" 10 = $2,000

PV="\frac{FV}{(1 + i)^{n} }=\frac{2,000}{(1 + 0.05)^{10} } = \frac{2,000}{1.62889} = \$1,227.83"